OpenStudy (anonymous):

I need help please, my brain is fried and I can't solve the simplest of equations............... x/3 - 7/8 = 2/5(x + 15/4) ........................ I am 3/4 of the way through but got stuck, can anyone help me?

OpenStudy (amistre64):

x/3 - 7/8 = 2/5(x + 15/4) ; *3 x - 21/8 = 6/5(x + 15/4) ; *8 8x - 21 = 48/5(x + 15/4) ; *5 40x - 105 = 48(x + 15/4)

OpenStudy (amistre64):

40x - 105 = 48(x + 15/4) *4 160x - 420 = 48(4x + 15)

OpenStudy (campbell_st):

multiply each term by the common denominator 120(x + 1/4) \[40(x + 15/4) - 105(x + 15/4) = 48\] expand, collect like terms then solve \[-65x - 243.75 = 48\] you can finish

OpenStudy (amistre64):

im assuming the (...) wasnt a denominator

OpenStudy (anonymous):

neither of those seem right, it's a simple equation that I need to solve for x

OpenStudy (amistre64):

well, we both got you to pretty much the end of it

OpenStudy (campbell_st):

good luck

OpenStudy (amistre64):

if the (..) is denominator; campbells good; if not, then mines good